OpenStudy (anonymous):

What operation is necessary to solve the equation that represents the statement: two point four is 8 multiplied by what number? Explain your answer.

OpenStudy (tainted):

Division, You can devide 8 by 2.4 to see what number it was multiplied by 8 / 2.4 = 3.33

OpenStudy (tainted):

divide*

OpenStudy (cwrw238):

8x = 2.4 divide both sides by 8

OpenStudy (cwrw238):

i believe its called the 'division property of equality' in the US

OpenStudy (cwrw238):

x = 2.4 / 8 = 0.3

OpenStudy (cwrw238):

im afraid tainted got it the wrong way round